# Sharp Capacity Scaling of Spectral Optimizers in Learning Associative Memory

### 27 Mar 2026 v1

@@ Abstract
Spectral optimizers such as Muon have recently shown strong empirical performance in large-scale language model training, but the source and extent of their advantage remain poorly understood. We study this question through the linear associative memory problem, a tractable model for factual recall in transformer-based models. In particular, we go beyond orthogonal embeddings and consider Gaussian inputs and outputs, which allows the number of stored associations to greatly exceed the embedding dimension. Our main result sharply characterizes the recovery rates of one step of Muon and SGD on the logistic regression loss under a power law frequency distribution. We show that the storage capacity of Muon significantly exceeds that of SGD, and moreover Muon saturates at a larger critical batch size. We further analyze the multi-step dynamics under a thresholded gradient approximation and show that Muon achieves a substantially faster initial recovery rate than SGD, while both methods eventually converge to the information-theoretic limit at comparable speeds. Experiments on synthetic tasks validate the predicted scaling laws. Our analysis provides a quantitative understanding of the signal amplification of Muon and lays the groundwork for establishing scaling laws across more practical language modeling tasks and optimizers.
@@ 1 Introduction
To investigate this question, we analyze the dynamics of Muon versus stochastic gradient descent (SGD) on the task of learning linear associative memory. The associative memory task, introduced in Cabannes et al. [6], Nichani et al. [38], provides a simple model of factual recall in language models [43, 1], and captures the ability of transformer-based models to store factual knowledge within the self-attention matrices. The goal is to store a collection of atomic associations (i.e., facts), expressed as $N$ pairs of input and output embeddings $\{(v_{i},u_{i})\}_{i\in[N]}\subset\mathbb{R}^{d}$, using a weight matrix ${\mathbf{W}}\in\mathbb{R}^{d\times d}$ so that $u_{i}\approx{\mathbf{W}}v_{i}$. We train $\mathbf{W}$} by casting this task as a multiclass logistic regression problem with logits given by $u_{j}^{\top}{\mathbf{W}}v_{i}$ and optimizing the cross-entropy loss. A formal description of the problem is given in Section 3.
@@
Theorem 1 (Informal version of Theorems 35).
Let $d$ be the embedding dimension and $B$ be the batch size, and suppose the $i$th item has power law frequency $p_{i}\propto i^{-\alpha}$ for $\alpha>1$. One step of Muon on the associative memory task recovers the top $\widetilde{\Theta}(\min\{d^{1+\frac{1}{2\alpha}},B^{\frac{1}{\alpha}}\})$ most frequent items, while one step of SGD recovers the top $\widetilde{\Theta}(\min\{d^{\frac{1}{2\alpha}},B^{\frac{1}{\alpha}}\})$ most frequent items.
This further implies that the critical batch size, beyond which increasing batch size does not yield performance gains, is much larger for Muon compared to SGD. The capacity exponents and batch size saturation predicted by our theory are empirically verified by our experiments (Figure 1).
@@
We also study the multi-step trajectories of Muon and SGD under a simplifying thresholded gradient update, and show the following scaling laws for the recovery rate (for brevity, we only state the population version below).
Theorem 2 (Informal version of Theorems 910).
@@
    Muon improves storage efficiency. In the population regime $B\rightarrow\infty$, one step of Muon is already able to recover the top $d^{1+\frac{1}{2\alpha}}$ items, while one step of SGD only recovers the top $d^{\frac{1}{2\alpha}}$ (see Figure 1(a)). Surprisingly, a single step of Muon is able to store more than $d$ items, which is the maximum when embeddings are constrained to be orthogonal. In other words, Muon is able to effectively store more features than dimensions via superposition.
@@
The rest of the paper is organized as follows. In Section 3, we formally define the associative memory task and the family of spectral optimizers considered in our analysis. Section 4 contains our main results on the scaling of a single step of Muon and SGD, as well as an argument that Muon is the asymptotically optimal one-step update. In Section 5, we extend our scaling analysis to multiple steps along the Muon and SGD trajectories. We conclude in Section 6 with simulations verifying our predicted scaling laws and batch size analysis, as well as experiments with transformers on a more sophisticated in-context recall task.
@@ 3 Setting: Associative Memory
[…]
@@
[…]
@@ 4 One Step of Muon
4.1 One-step recovery of Muon and SGD
@@
Theorem 3 (one-step recovery of Muon).
Let $u_{i},v_{i}$ for $i\in[N]$ be i.i.d. $\mathcal{N}(0,\frac{1}{d}{\mathbf{I}}_{d})$ vectors. Let ${\mathbf{G}}_{0}=-\nabla_{{\mathbf{W}}}L({\mathbf{W}}_{0};\mathcal{B})$ be the negative gradient at initialization of the empirical loss on a minibatch $\mathcal{B}$ of size $B$, or the population loss (equivalently $B=\infty$). Suppose $N=\operatorname{poly}(d)$, $N\gtrsim d^{2\alpha+1}$ and set
@@
Then with high probability, the one-step Muon update ${\mathbf{W}}_{1}\propto h_{\lambda}({\mathbf{G}}_{0})$ recovers all items up to
$i\lesssim\min\left\{i^{\star},B^{\frac{1}{\alpha}}(\log d)^{-\frac{1}{\alpha}}\right\},\quad i^{\star}\asymp d^{1+\frac{1}{2\alpha}}(\log d)^{-2-\frac{5}{\alpha}}.$
@@
and this allows us to recover $\widetilde{\Theta}(d^{1+\frac{1}{2\alpha}})$ items. We also note that the condition $N\gtrsim d^{2\alpha+1}$ can be removed by considering correlation loss.
@@
Corollary 4.
Taking the learning rate $\eta\asymp(\log d)^{-4}\sqrt{d}$ in the setting of Theorem 3, the one-step update ${\mathbf{W}}_{1}=\eta h_{\lambda}({\mathbf{G}}_{0})$ achieves loss
$L({\mathbf{W}}_{1})\leq\widetilde{O}\quantity(\max\left\{d^{\frac{1}{2\alpha}+\frac{1}{2}-\alpha},B^{\frac{1}{\alpha}-1}\right\})$
@@
The proof of Theorem 3 is developed throughout Appendices AB. A sketch of the main ideas is provided in Section 4.2. We start by observing from Eq. (2) that the gradient of the cross-entropy loss $L({\mathbf{W}})$ at initialization is roughly ${\mathbf{G}}_{0}\approx\sum_{i}p_{i}u_{i}v_{i}^{\top}$. If the embeddings $u_{i},v_{i}$ were orthogonal, Muon would then output ${\mathbf{W}}_{1}\propto h_{\lambda}({\mathbf{G}}_{0})\approx\sum_{i}u_{i}v_{i}^{\top}$ which already classifies all items correctly; however this constrains the number of items $N\leq d$, far less than the information-theoretic optimum $\widetilde{\Theta}(d^{2})$. In our non-orthogonal setting, $N$ can be much larger, but we must now account for the correlations between embeddings.
Let us now focus on each signal term $u_{i}^{\top}{\mathbf{W}}_{1}v_{i}$. The main contribution comes from the aligned rank-one spike $p_{i}u_{i}v_{i}^{\top}$ in the gradient ${\mathbf{G}}_{0}$. At the same time, a large fraction of the singular values of ${\mathbf{G}}_{0}$ lie below $d^{-\alpha}$. The map $h_{\lambda}$ amplifies these by a factor of $\lambda^{-1}\sim d^{\alpha}$, which also boosts the spike magnitude $p_{i}$ by the same factor, allowing us to recover items with lower frequencies. Thus, $\lambda$ effectively acts as a scale of resolution for the singular spectrum; the implications of this is discussed in Section 4.3.
In contrast, we show the following bound on the number of recovered items for vanilla gradient descent on the same objective.
Theorem 5 (one-step recovery of SGD).
In the setting of Theorem 3 and $N\gtrsim d$, the number of items recovered by the one-step SGD update ${\mathbf{W}}_{1}^{\operatorname{SGD}}=\eta{\mathbf{G}}_{0}$ is
@@
4.2 Proof sketch of Theorem 3
We start by approximating the gradient as a sum of independent rank-one terms:
@@ 4.2 Proof sketch of Theorem 3
and defining the logits $\gamma_{ij}:=u_{j}^{\top}h_{\lambda}({\mathbf{G}})v_{i}$ (omitting $\eta$). We analyze the signal and interaction terms separately.
@@
By definition, $\gamma_{ij}=\phi(q_{i})$, which we analyze via Taylor expansion. The slope $\phi^{\prime}(0)$ can be computed explicitly via the Daleckii–Krein formula (Proposition 11), which computes the Fréchet derivative of a matrix function along a specified perturbation direction. Let ${\mathbf{G}}_{-i}={\mathbf{A}}{\mathbf{S}}{\mathbf{B}}^{\top}$ be the SVD with singular values ${\mathbf{S}}=\operatorname{diag}(s_{1},\cdots,s_{d})$ in decreasing order and $a={\mathbf{A}}^{\top}u_{i}$, $b={\mathbf{B}}^{\top}v_{i}$. Then $\phi^{\prime}(0)$ is given as the sum of nonnegative terms
@@
We focus on the first term. As $a,b$ are i.i.d. Gaussian conditioned on ${\mathbf{G}}_{-i}$, $(a_{k}b_{\ell}-a_{\ell}b_{k})^{2}\approx 1/d^{2}$. Since $z\mapsto h_{\lambda}(z)/z$ is decreasing, the sum is dominated by small singular values. We then show that the ‘bulk’ singular value $s_{d/2}=\widetilde{O}(d^{-\alpha})$ with high probability (Lemmas 14, 17). Hence choosing $\lambda$ at the same scale,
@@
The interaction logits turn out to be much more challenging to control. A naive approach is to use that $h_{\lambda}$ is $\lambda^{-1}$-Lipschitz with respect to operator norm (Proposition 26), so $(u_{j},v_{j})\mapsto\gamma_{ij}=u_{j}^{\top}h_{\lambda}({\mathbf{G}}_{-j}+q_{j}u_{j}v_{j}^{\top})v_{i}$ is a Lipschitz mapping of Gaussians and so exhibits good concentration. This argument works when either $q_{i}$ or $q_{j}\ll\lambda$, but fails to bound ‘large’ interactions where both $i,j\leq r$ for a threshold $r\approx d$. For these terms, we first invoke a block resolvent integral representation amenable to series expansion, then develop a nonasymptotic perturbative analysis reminiscent of moment methods in random matrix theory. A technical overview is provided in Appendix B.1 for the interested reader. In the end, we show:

### 28 Apr 2026 v2

@@ Abstract
Spectral optimizers such as Muon have recently shown strong empirical performance in large-scale language model training, but the source and extent of their advantage remain poorly understood. We study this question through the linear associative memory problem, a tractable model for factual recall in transformer-based models. In particular, we go beyond orthogonal embeddings and consider Gaussian inputs and outputs, which allows the number of stored associations to greatly exceed the embedding dimension. Our main result sharply characterizes the recovery rates of one step of Muon, SGD, and Newton’s method on the logistic regression loss under a power law frequency distribution. We show that the storage capacity of Muon significantly exceeds that of SGD, and even matches Newton’s method while only using first-order information. Moreover, Muon saturates at a larger critical batch size. We further analyze the multi-step dynamics under a thresholded gradient approximation and show that Muon achieves a substantially faster initial recovery rate than SGD, while both methods eventually converge to the information-theoretic limit at comparable speeds. Experiments on synthetic tasks validate the predicted scaling laws. Our analysis provides a quantitative understanding of the signal amplification of spectral preconditioners and lays the groundwork for establishing scaling laws across more practical language modeling tasks and optimizers.
@@ 1 Introduction
To investigate this question, we analyze the dynamics of Muon versus stochastic gradient descent (SGD) and Newton’s method on the task of learning linear associative memory. The associative memory task, introduced in Cabannes et al. [6], Nichani et al. [39], provides a simple model of factual recall in language models [44, 1], and captures the ability of transformer-based models to store factual knowledge within the self-attention matrices. The goal is to store a collection of atomic associations (i.e., facts), expressed as $N$ pairs of input and output embeddings $\{(v_{i},u_{i})\}_{i\in[N]}\subset\mathbb{R}^{d}$, using a weight matrix ${\mathbf{W}}\in\mathbb{R}^{d\times d}$ so that $u_{i}\approx{\mathbf{W}}v_{i}$. We train $\mathbf{W}$} by casting this task as a multiclass logistic regression problem with logits given by $u_{j}^{\top}{\mathbf{W}}v_{i}$ and optimizing the cross-entropy loss. A formal description of the problem is given in Section 3.
@@
Theorem 1.1 (Informal version of Theorems 4.14.34.4).
Let $d$ be the embedding dimension and $B$ be the batch size, and suppose the $i$th item has power law frequency $p_{i}\propto i^{-\alpha}$ for $\alpha>1$. One step of Muon on the associative memory task recovers the top $\widetilde{\Theta}(\min\{d^{1+\frac{1}{2\alpha}},B^{\frac{1}{\alpha}}\})$ most frequent items, matching the Newton update, while one step of SGD recovers the $\widetilde{\Theta}(\min\{d^{\frac{1}{2\alpha}},B^{\frac{1}{\alpha}}\})$ items.
Surprisingly, Muon is even able to match Newton’s method – the gold standard of curvature-aware optimization – using only first-order information, demonstrating the power of spectral preconditioning. The theorem also implies that the critical batch size, beyond which increasing batch size does not yield performance gains, is much larger for Muon compared to SGD. The capacity exponents and batch size saturation predicted by our theory are empirically verified by our experiments (Figure 1).
@@
Furthermore, we study the multi-step trajectories of Muon and SGD under a simplifying thresholded gradient update, and show the following scaling laws for the recovery rate (for brevity, we only state the population version below).
Theorem 1.2 (Informal version of Theorems 5.45.5).
@@
    Muon improves storage efficiency. In the population regime $B\rightarrow\infty$, one step of Muon is able to recover the top $d^{1+\frac{1}{2\alpha}}$ items, matching Newton’s method, while one step of SGD only recovers the top $d^{\frac{1}{2\alpha}}$ (see Figure 1(a)). Noticeably, a single step of Muon is able to store more than $d$ items, which is the maximal value when embeddings are constrained to be orthogonal. In other words, Muon effectively stores more features than dimensions via superposition.
@@
The rest of the paper is organized as follows. In Section 3, we formally define the associative memory task and the family of spectral optimizers considered in our analysis. Section 4 contains our main results on the scaling of a single step of Muon, SGD, and Newton. In Section 5, we argue that Muon is the asymptotically optimal one-step update, and extend our scaling analysis to multiple steps along the Muon and SGD trajectories. We conclude in Section 6 with simulations verifying our predicted scaling laws and batch size analysis, as well as experiments with transformers on a more sophisticated in-context recall task.
@@ 3 Setting: Associative Memory
$L({\mathbf{W}}):=\mathbb{E}_{i\sim p}[-\log p_{{\mathbf{W}}}(i\mid i)]=-\sum_{i\in[N]}p_{i}\Bigg(u_{i}^{\top}{\mathbf{W}}v_{i}-\log\sum_{j\in[N]}\exp(u_{j}^{\top}{\mathbf{W}}v_{i})\Bigg).$ (1)
@@
$L({\mathbf{W}};\mathcal{B}):=\frac{1}{B}\sum_{i\in\mathcal{B}}-\log\hat{p}_{{\mathbf{W}}}(i\mid i)=-\sum_{i\in[N]}q_{i}\Bigg(u_{i}^{\top}{\mathbf{W}}v_{i}-\log\sum_{j\in[N]}\exp(u_{j}^{\top}{\mathbf{W}}v_{i})\Bigg),$
@@ 4 One Step of Muon
4.1 One-step recovery of Muon
@@
Theorem 4.1 (one-step recovery of Muon).
Let $u_{i},v_{i}$ for $i\in[N]$ be i.i.d. $\mathcal{N}(0,\frac{1}{d}{\mathbf{I}}_{d})$ vectors. Let ${\mathbf{G}}_{0}=-\nabla_{{\mathbf{W}}}L({\mathbf{W}}_{0};\mathcal{B})$ be the negative gradient at initialization of the empirical loss on a minibatch $\mathcal{B}$ of size $B$, or the population loss (equivalently $B=\infty$). Suppose $N=\operatorname{poly}(d)$, $N\gtrsim d^{2\alpha+2}$ and set
@@
Then with high probability, the one-step Muon update ${\mathbf{W}}_{1}^{\operatorname{Muon}}\propto h_{\lambda}({\mathbf{G}}_{0})$ recovers all items up to
$i\lesssim\min\left\{i^{\star},B^{\frac{1}{\alpha}}(\log d)^{-\frac{1}{\alpha}}\right\},\quad i^{\star}\asymp d^{1+\frac{1}{2\alpha}}(\log d)^{-2-\frac{5}{\alpha}}.$ (5)
@@
and this allows us to recover $\widetilde{\Theta}(d^{1+\frac{1}{2\alpha}})$ items. We also note that the condition $N\gtrsim d^{2\alpha+2}$ can be removed by considering correlation loss.
@@
Corollary 4.2.
Taking the learning rate $\eta\asymp(\log d)^{-4}\sqrt{d}$ in the setting of Theorem 4.1, the one-step update ${\mathbf{W}}_{1}^{\operatorname{Muon}}=\eta h_{\lambda}({\mathbf{G}}_{0})$ achieves loss
$L({\mathbf{W}}_{1}^{\operatorname{Muon}})\leq\widetilde{O}\quantity(\max\left\{d^{\frac{1}{2\alpha}+\frac{1}{2}-\alpha},B^{\frac{1}{\alpha}-1}\right\})$ (6)
@@
The proof of Theorem 4.1 is developed throughout Appendices AB; a sketch of the main ideas is provided in Section 4.3. Here, we make some basic observations. From Eq. (2), the gradient of the cross-entropy loss $L({\mathbf{W}})$ at initialization is roughly ${\mathbf{G}}_{0}\approx\sum_{i}p_{i}u_{i}v_{i}^{\top}$. If the embeddings $u_{i},v_{i}$ were orthogonal, Muon would then output ${\mathbf{W}}_{1}^{\operatorname{Muon}}\propto h_{\lambda}({\mathbf{G}}_{0})\approx\sum_{i}u_{i}v_{i}^{\top}$ which already classifies all items correctly; however this constrains the number of items $N\leq d$, far less than the information-theoretic optimum $\widetilde{\Theta}(d^{2})$. In our non-orthogonal setting, $N$ can be much larger, but we must now account for the correlations between embeddings.
Let us now focus on each signal term $u_{i}^{\top}h_{\lambda}({\mathbf{G}}_{0})v_{i}$. The main contribution comes from the aligned rank-one spike $p_{i}u_{i}v_{i}^{\top}$ in the gradient ${\mathbf{G}}_{0}$. We quantify this through an add-back-in argument: starting from the leave-one-out component ${\mathbf{G}}_{-i}=\sum_{j\neq i}p_{j}u_{j}v_{j}^{\top}$, we analyze how quickly the logit grows as the $u_{i}v_{i}^{\top}$ spike coefficient increases from 0 to $p_{i}$. At the same time, a large fraction of the singular values of ${\mathbf{G}}_{0}$ lie below $d^{-\alpha}$. The map $h_{\lambda}$ amplifies these by a factor of $\lambda^{-1}\sim d^{\alpha}$, which also boosts the logit growth rate by the same factor, allowing us to recover items with lower frequencies $p_{i}$. Thus, $\lambda$ effectively acts as a scale of resolution for the singular spectrum; the implications of this is discussed in Section 5.1.
4.2 One-step recovery of SGD and Newton
Stochastic gradient descent.
In contrast with Theorem 4.1, we next prove a tight bound on the number of recovered items for vanilla SGD on the same objective.
Theorem 4.3 (one-step recovery of SGD).
In the setting of Theorem 4.1 and $N\gtrsim d$, the number of items recovered by the one-step SGD update ${\mathbf{W}}_{1}^{\operatorname{SGD}}=\eta{\mathbf{G}}_{0}$ is
@@ 4.2 One-step recovery of SGD and Newton
Newton’s method.
To put the capacity gain of Muon into perspective, we now show that for the first optimization step, Muon in fact matches the recovery rate of Newton’s method for linear associative memory. The Newton update is defined as the direction towards the minimizer of the local quadratic approximation to $L$:
${\mathbf{W}}\leftarrow{\mathbf{W}}+\eta\cdot[\nabla_{{\mathbf{W}}}^{2}L({\mathbf{W}};\mathcal{B})]^{-1}{\mathbf{G}}.$
Theorem 4.4 (one-step recovery of Newton).
Denote the Hessian of the loss at initialization as $\mathcal{H}=\nabla_{{\mathbf{W}}}^{2}L({\mathbf{W}}_{0};\mathcal{B})$. In the setting of Theorem 4.1, if $B=\widetilde{\Omega}(d^{\alpha})$ and $\eta=\widetilde{\Theta}(1/\sqrt{d})$, the one-step Newton update ${\mathbf{W}}_{1}^{\operatorname{Newton}}=\eta\mathcal{H}^{-1}[{\mathbf{G}}_{0}]$ achieves the same recovery rate Eq. (5) and loss decrease Eq. (6) of Muon, up to log factors.
This result is particularly surprising because Newton’s method is often considered the gold standard of local second-order or curvature-aware optimization. Many popular preconditioned optimizers – such as Gauss-Newton, Adagrad [10], K-FAC [36], and Shampoo [17] – can be viewed as tractable approximations which avoid the cost of computing and inverting the full Hessian. Theorem 4.4 shows that Muon can match Newton without accessing any second-order information, even though the updates are structurally different.
To gain intuition on this comparison, we explicitly write down the Newton update and compare with Muon (taking $\lambda=0$ for simplicity):
${\mathbf{W}}_{1}^{\operatorname{Newton}}\propto\Big(\underbrace{\textstyle\frac{1}{N}\sum_{i}u_{i}u_{i}^{\top}-\bar{u}\bar{u}^{\top}}_{=:\boldsymbol{\Sigma}_{u}}\Big)^{-1}{\mathbf{G}}_{0}\Big(\underbrace{\textstyle\sum_{i}q_{i}v_{i}v_{i}^{\top}}_{=:{\mathbf{M}}_{v}}\Big)^{-1}\quad\text{vs.}\quad{\mathbf{W}}_{1}^{\operatorname{Muon}}\propto{\mathbf{G}}_{0}({\mathbf{G}}_{0}^{\top}{\mathbf{G}}_{0})^{-1/2}.$ (7)
The Hessian admits a Kronecker factorization $\mathcal{H}={\mathbf{M}}_{v}\otimes\boldsymbol{\Sigma}_{u}$ at initialization, thus the Newton update is equivalent to the K-FAC update with preconditioning on both sides, while Muon is only preconditioned on the right. In particular, the left preconditioner $\boldsymbol{\Sigma}_{u}$ is the (unweighted) empirical covariance matrix, which has the effect of whitening the unembedding vectors $u_{i}$. We note that the batch size condition $B\gtrsim d^{\alpha}$ is necessary for the Hessian to be well-conditioned (Lemma C.3). Regarding the proof, we analyze the Newton step by applying the add-back-in argument to each spike in ${\mathbf{G}}_{0},{\mathbf{M}}_{v}$. Compared to Muon, the analysis is relatively straightforward as we can directly apply the Sherman–Morrison formula to ${\mathbf{M}}_{v}^{-1}$ to compute the change in logits.
Handling anisotropic embeddings.
While Muon matches Newton’s method in our isotropic Gaussian setting, an important limitation appears for anisotropic data. Suppose the unembedding and embedding vectors follow $u_{i}\sim\mathcal{N}(0,\boldsymbol{\Xi}_{u})$ and $v_{i}\sim\mathcal{N}(0,\boldsymbol{\Xi}_{v})$ with general covariance matrices $\boldsymbol{\Xi}_{u},\boldsymbol{\Xi}_{v}$. It is apparent from Eq. (7) that the logits $u_{j}^{\top}{\mathbf{W}}_{1}^{\operatorname{Newton}}v_{i}$ (and thus the estimated likelihoods) of the Newton update are invariant under the transformations $u_{i}\mapsto\boldsymbol{\Xi}_{u}^{-1/2}u_{i}$ and $v_{i}\mapsto\boldsymbol{\Xi}_{v}^{-1/2}v_{i}$, and therefore achieve the same recovery rate as in Theorem 4.4. By contrast, the polar map is not invariant under these transformations, so Muon cannot in general be expected to retain the same rate. This is consistent with the experiments in Figure 6: when $u_{i}$ and $v_{i}$ have identity covariance, Muon achieves capacity comparable to Newton’s method, but the gap widens as the unembedding vectors become more anisotropic.
4.3 Proof sketch of Theorem 4.1
To start, we approximate the gradient as a sum of independent rank-one terms:
@@
and define the logits $\gamma_{ij}:=u_{j}^{\top}h_{\lambda}({\mathbf{G}})v_{i}$ (omitting $\eta$). We analyze the signal and interaction terms separately.
@@
By definition, $\gamma_{ij}=\phi(q_{i})$, which we analyze via Taylor expansion. The slope $\phi^{\prime}(0)$ can be computed explicitly via the Daleckii–Krein formula (Proposition A.1), which computes the Fréchet derivative of a matrix function along a specified perturbation direction. Let ${\mathbf{G}}_{-i}={\mathbf{A}}{\mathbf{S}}{\mathbf{B}}^{\top}$ be the SVD with singular values ${\mathbf{S}}=\operatorname{diag}(s_{1},\cdots,s_{d})$ in decreasing order and $a={\mathbf{A}}^{\top}u_{i}$, $b={\mathbf{B}}^{\top}v_{i}$. Then $\phi^{\prime}(0)$ is given as the sum of nonnegative terms
@@
We focus on the first term. As $a,b$ are i.i.d. Gaussian conditioned on ${\mathbf{G}}_{-i}$, $(a_{k}b_{\ell}-a_{\ell}b_{k})^{2}\approx 1/d^{2}$. Since $z\mapsto h_{\lambda}(z)/z$ is decreasing, the sum is dominated by small singular values. We then show that the ‘bulk’ singular value $s_{d/2}=\widetilde{O}(d^{-\alpha})$ with high probability (Lemmas A.4, A.7). Hence choosing $\lambda$ at the same scale,
@@
The interaction logits turn out to be much more challenging to control. A naive approach is to use that $h_{\lambda}$ is $\lambda^{-1}$-Lipschitz w.r.t. operator norm (Proposition B.9), so $(u_{j},v_{j})\mapsto\gamma_{ij}=u_{j}^{\top}h_{\lambda}({\mathbf{G}}_{-j}+q_{j}u_{j}v_{j}^{\top})v_{i}$ is a Lipschitz mapping of Gaussians and so exhibits good concentration. This argument works when either $q_{i}$ or $q_{j}\ll\lambda$, but fails to bound ‘large’ interactions where both $i,j\leq r$ for a threshold $r\approx d$. For these terms, we first invoke a block resolvent integral representation amenable to series expansion, then develop a nonasymptotic perturbative analysis reminiscent of moment methods in random matrix theory. A technical overview is provided in Appendix B.1 for the interested reader. In the end, we show: